# Elimination Distance to Dominated Clusters

### v1

@@ 4 Algorithm on unbreakable graphs based on Annotated Partial Domination
As sketched above, we first solve Dominated Cluster Deletion and Elimination Distance to Dominated Clusters on unbreakable graphs. Our algorithm runs efficiently on all graph classes for which we can efficiently solve the Annotated Partial Domination problem, that is, for classes on which a parameter $t(\cdot)$ satisfying 1 is bounded.
In the following, $q,k$, and $d$ will always be non-negative integers, and we will not always quantify them for readability. We will also tacitly assume that our graphs have at least $12q+$ 1 vertices. Then, whenever we delete a set $S$ of at most $k$ vertices there will be a unique connected component $C_{0}$ of $G-S$ with more than $q$ vertices that we call the large connected component. When a graph has less vertices, we can solve all our problems by brute-force.
@@
Let $G$ be a $(q,k)$-unbreakable graph, and let $v\in V(G)$. Then $G-v$ is $(q,k-1)$-unbreakable.
@@
Assume that there is a connected component of $G-v$ that is not $(q,k-1)$-unbreakable. Then, there is a separation $(A,B)$ of $G-v$ such that $|A|\geq q$, $|B|\geq q$ and $|A\cap B|\leq k-1$. By adding the vertex $v$ to the separator, we obtain a separation $(A^{\prime},B^{\prime})$ of $G$ such that $|A^{\prime}|>q$, $|B^{\prime}|>q$ and $|A^{\prime}\cap B^{\prime}|\leq k$ contracting that $G$ is $(q,k)$-unbreakable.
@@
Let $S^{\prime}$ with $|S^{\prime}|\leq k$ be a solution for Dominated Cluster Deletion with parameters $d$ and $k$. Let $A=S^{\prime}\cup C_{0}$, where $C_{0}$ is the large connected component of $G-S^{\prime}$ and $B=S^{\prime}\cup\bigcup_{1\leq j\leq m}C_{j}$, where $C_{1},\ldots,C_{m}$ are the other connected components of $G-S^{\prime}$. Then $(A,B)$ is a separation of order at most $k$, hence $|B|\leq q$, as $G$ is $(q,k)$-unbreakable. As $G$ is a positive instance of Dominated Cluster Deletion, $C_{0}$ can be dominated by a set $D_{0}$ of size at most $d$. Then $D_{0}\cup B$ is dominating set of $G$ of size at most $q+d$.
@@ 4.1 Dominated Cluster Deletion
Note that $S$ may be the empty set in the third item above if $G-S^{\prime}$ has only one connected component. Figure 3 shows examples of skeletons for dominated cluster deletion.
@@
Assume that the Case 1 is not satisfied. Meaning that no vertices from the skeleton are in the dominating set $X$. Denote by $C^{\prime}_{0}$ the large connected component of $G-S^{\prime}$ for a set $S^{\prime}\supseteq S$ witnessing that $S$ is a skeleton, hence, $S$ contains exactly the vertices of $S^{\prime}$ that have a neighbor in $C^{\prime}_{0}$ and a second neighbor in $G-N[C^{\prime}_{0}]$. Let $C_{0}$ be the large connected component of $G-S$. Note that it contains $C^{\prime}_{0}$.
Let $A=S\cup C_{0}$ and $B=S\cup\bigcup_{1\leq j\leq m}C_{j}$, where $C_{1},\ldots,C_{m}$ are the other connected components of $G-S$. Then $(A,B)$ is a separation of order at most $|S|\leq k$, hence $|B|\leq q$, as $G$ is $(q,k)$-unbreakable. Pick an arbitrary element $v\in S$. By definition of $S$, $v$ has a neighbor in $C_{0}$ and in a second connected component, say $C_{1}$ of $G-S$. This neighbor must be a vertex $w$ of degree at most $q$ as its neighborhood lies in $G-C_{0}\subseteq B$. If $w$ is in $X$, then this vertex $v$ satisfies Case 2. Otherwise, $w$ must have a neighbor $x$ in $X$ (since $X$ is a dominating set). Since we are not in Case 1, this $x$ is not in $S$, and is therefore in $C_{1}$. Hence, similarly to $w$, we have that $x$ has degree at most $q$, and we are in Case 3.
Lemma 4.6 gives rise to a bounded search tree branching algorithm that branches over the possible choices for a vertex from $S$ given by the conditions of the lemma.
@@
The skeleton-algorithm with parameter $q,k,d$ over a graph $G$ returns a family of sets of vertices, each set of size at most $k$. First, it computes a dominating set $X$ of size at most $q+d$ (returning the empty family if there is no such set). Second, it computes the set $X\cup Y\cup Z$, where $Y$ is the set of all vertices in the neighborhood $N(x)$ of a vertex $x\in X$ with $\deg(x)\leq q$, and the set $Z$ is the set of all vertices in the neighborhood $N(y)$ of a vertex $y\in Y$ where $\deg(y)\leq q$. Third, call the algorithm with parameters $q,k-1,d$ on $G-v$ for every $v$ in $X\cup Y\cup Z$ (and get a family $\mathcal{F}_{v}$). Finally, for every $v$ in $X\cup Y\cup Z$ add $v$ to every set in $\mathcal{F}_{v}$, and return the empty set and the union of all the families $\mathcal{F}_{v}$.
Note that in the above definition, if a family $\mathcal{F}_{v}$ is empty, adding $v$ to every set still results in an empty family (e.g. in the case where $G-v$ does not have a $q+d$ dominating set). To not be confused with $\mathcal{F}_{v}$ having the empty set as one of the family member (e.g. for branches of the algorithm that stops before reaching $k=0$).
@@
We now show that the skeleton algorithm 1) runs efficiently and outputs a small family of sets. This is proved in Lemma 4.9, and 2) if the given graph was a positive instance to Dominated Cluster Deletion, then the skeleton algorithm outputs all possible skeletons; see Lemma 4.11.
@@
Let $t(\cdot)$ be a graph parameter satisfying 1, then the skeleton-algorithm runs in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$, and outputs a family of at most $g(q,k,d)$ sets where $f$ and $g$ are computable functions.
@@
Computing $X,Y$ and $Z$ takes time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$ thanks to the fact that $t(\cdot)$ satisfies 1. We then branch over $(q+d)q^{2}$ many choices for $v$, and recurse at most $k$ times.
@@
By induction on $k$. If $k=0$, we simply output the empty set. Otherwise, by Lemma 4.3 there is such a set $X$, and by Lemma 4.6 at least one element $v$ of $X$ is in $S$. For the correct choices of $v$, $G-v$ is a positive instance for Dominated Cluster Deletion with parameter $k-1,d$, and by Lemma 4.1, $G-v$ is $(q,k-1)$-unbreakable, so the recursive call to Skeleton produces the expected sets by induction.
@@
Given a $(q,k)$-unbreakable graph $G$ and a set $S$ of at most $k$ vertices, we can test in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$, whether $S$ is a skeleton for Dominated Cluster Deletion with parameters $k$ and $d$, and $t(\cdot)$ satisfies 1.
@@
We need to check the existence of a set $S^{\prime}\supseteq S$ such that:
@@
Note that by these conditions the large connected component $C_{0}$ of $G-S$ is separated from the small connected components by $S$. What remains to be done is hence the following.
@@
    Find an optimal solution for $G-S-C_{0}$. This can be done in time $f(k,d)$, because this subgraph has size at most $q$ and yields a value $k^{\prime\prime}$ of vertices that need to be deleted. If $k^{\prime\prime}>k-k^{\prime}$ we reject the instance (in this branch).
@@
    Test whether an additional set $W$ of $k^{\prime\prime\prime}=k-k^{\prime}-k^{\prime\prime}$ vertices can be deleted from $C_{0}$ such that $G[C_{0}-W]$ can be dominated by $d$ vertices. Note that by assumption on $S$ and $S^{\prime}$ the deletion of $W$ will not cause $C_{0}$ to break into multiple further connected components. Hence, the remaining problem is an instance of Partial Domination with parameter $k^{\prime\prime\prime}$. If $S$ is a skeleton, such a set exists (as $S^{\prime}\cap C_{0}$ is a candidate) and hence, if $G$ is a positive instance of Dominated Cluster Deletion a positive solution will be found. Since $t(\cdot)$ is a graph parameter satisfying 1, we can test the existence of such $W$ in time $f(k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$. If no solution is found we reject the instance (in this branch).
The combination of Lemmas 4.6, 4.9, 4.11 and 4.13 yields the following theorem.
@@
Dominated Cluster Deletion on $(q,k)$-unbreakable graphs can be solved in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$, where $t(\cdot)$ is a parameter satisfying 1.
@@ 4.2 Elimination Distance to Dominated Clusters
First let’s properly define Elimination Distance to Dominated Clusters.
@@
  • $G$ is connected and there is a vertex $v$ such that $G-v$ has elimination distance $k-1$ to $\mathscr{C}$, or
@@
In our case we measure, for every $d$, the elimination distance to the class $\mathscr{C}_{d}$ of $d$-dominated clusters i.e., graphs where every connected component has a $d$-dominating set. We interchangeably write “elimination distance $k$ to $d$-dominated cluster”, and “a positive instance for $(k,d)$-Elimination Distance to Dominated Clusters”. For algorithmic purposes, the above definition is not the most useful. We will instead look at the set of vertices that needs to be deleted to get $d$-dominated clusters. While this set is of unbounded size, it has a nice tree structure that is much more useful for our algorithms. This characterization is pretty standard, see the definition of elimination order [9, Proposition 4.3]
@@
Let $G$ be a graph and $S\subseteq V(G)$. We say that $S$ is tree-structured if there is a rooted tree $T$ and a mapping $\lambda:S\rightarrow V(T)$ such that every path in $G$ between two vertices $u,v$ of $S$ must contain a common ancestor in $T$ of $\lambda(u)$ and $\lambda(v)$. The tree $T$ is called an elimination tree for $S$. The elimination depth of a tree-structured set with respect to the tree $T$ is the depth of $T$. The elimination depth of $S$ is the minimum elimination depth of $S$ over all elimination trees for $S$.
@@
A graph $G$ is a positive instance for $(k,d)$-Elimination Distance to Dominated Clusters if and only if there is a tree-structured set $S$ with elimination depth $k$ such that every connected component of $G-S$ has a $d$-dominating set.
@@
Let $G$ be a $(q,k)$-unbreakable graph. Let $S$ be a tree-structured set with an elimination tree $T$ of depth at most $k$. Then
@@
    the largest connected component $C_{0}$ of $G-S$ is uniquely determined, and
@@
    for every component $C$ of $G-S$ there is a subset $S(C)\subseteq S$ of size at most $k$ that separates $C$ from the rest, that is, $G-S(C)$ has $C$ as one of its components.
@@
Let $C$ be a connected component of $G-S$. As $S$ is tree-structured by $T$, $C$ is not adjacent to two $\preccurlyeq_{T}$-incomparable elements, that is, the set $S(C)$ of neighbors of $C$ in $S$ consists of $\preccurlyeq_{T}$-comparable elements. $C$ is a connected component both of $G-S$ and of $G-S(C)$, as all neighbors of $C$ in $S$ are collected in $S(C)$.
Assume now that there are two connected components $C_{0},C_{1}$ of $G-S$ of size greater than $q$. However, $C_{0}$ and $C_{1}$ are separated by $S(C_{0})$ of size at most $k$, which contradicts the $(q,k)$-unbreakability of $G$ (assuming that $G$ has at least $12q+$ 1 vertices). Hence, the largest connected component $C_{0}$ of $G-S$ is the unique connected component of size greater than $q$.
Lemma 4.21 (Analog of Lemma 4.3).
@@
The proof can be taken verbatim from Lemma 4.3 with $S^{\prime}$ replaced by $S^{\prime}(C_{0})$.
@@
The following observation follows immediately from Lemma 4.19. {observation} Every skeleton has size at most $k$.
Figure 4: Example of a skeleton for elimination distance to dominated clusters for parameter $k=3$ and $d=1$ on a (12,3)-unbreakable graph (left). The skeleton $S$ is in red, the witness $S^{\prime}$ also contains the crossed-out vertices. A tree structure for $S^{\prime}$ of elimination depth $3$ is on the right.
Now the analog of Lemma 4.6 also holds for skeletons for elimination distance to dominated clusters. The proof can be taken verbatim with $S^{\prime}$ replaced by $S^{\prime}(C_{0})$.
@@
Hence, the algorithm to guess a skeleton $S$ can be carried out in exactly the same way as for Dominated Cluster Deletion. We then guess a linear order on $S$ (which takes time at most $\mathcal{O}(k!)$). In the final step when we aim to extend a fixed skeleton $S$ to a tree-structured solution $S^{\prime}$, the only difference is that we have to solve the base problem Elimination Distance to Dominated Clusters for the small part $G-S-C^{\prime}_{0}$, which again can be simply solved by brute-force. Note that because $G-S-C^{\prime}_{0}$ is small we can also find an elimination tree $T$ for $S^{\prime}$ and embed $S$ according to the originally guessed order. We conclude the main theorem of this section. The full algorithm is presented in the appendix.
@@
Elimination Distance to Dominated Clusters on $(q,k)$-unbreakable graphs can be solved in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$, where $t(\cdot)$ is a parameter satisfying 1.

### v2

@@ 4 Algorithm on unbreakable graphs based on Annotated Partial Domination
As sketched above, we first solve Dominated Cluster Deletion and Elimination Distance to Dominated Clusters on unbreakable graphs. Our algorithm runs efficiently on all graph classes for which we can efficiently solve the Annotated Partial Domination problem, that is, for classes on which a parameter $t(\cdot)$ satisfying ˜1 is bounded.
In the following, $q,k$, and $d$ will always be non-negative integers, and we will not always quantify them for readability. We will also tacitly assume that our graphs have at least $2q+1$ vertices. Then, whenever we delete a set $S$ of at most $k$ vertices there will be a unique connected component $C_{0}$ of GSG-S with more than $q$ vertices that we call the large connected component. When a graph has less vertices, we can solve all our problems by brute-force.
@@
Let $G$ be a $(q,k)$-unbreakable graph, and let $v\in V(G)$. Then GvG-v is (q,k1)(q,k-1)-unbreakable.
@@
Assume that GvG-v that is not (q,k1)(q,k-1)-unbreakable. Then, there is a separation $(A,B)$ of GvG-v such that $|A|\geq q$, $|B|\geq q$ and $|A\cap B|\leq k-1$. By adding the vertex $v$ to the separator, we obtain a separation $(A^{\prime},B^{\prime})$ of $G$ such that $|A^{\prime}|>q$, $|B^{\prime}|>q$ and $|A^{\prime}\cap B^{\prime}|\leq k$ contradicting that $G$ is $(q,k)$-unbreakable.
@@
Let $S^{\prime}$ with $|S^{\prime}|\leq k$ be a solution for Dominated Cluster Deletion with parameters $d$ and $k$. Let $A=S^{\prime}\cup C_{0}$, where $C_{0}$ is the large connected component of $G-S^{\prime}$ and $B=S^{\prime}\cup\bigcup_{1\leq j\leq m}C_{j}$, where $C_{1},\ldots,C_{m}$ are the other connected components of $G-S^{\prime}$. Then, $(A,B)$ is a separation of order at most $k$, hence $|B|\leq q$, as $G$ is $(q,k)$-unbreakable. As $G$ is a positive instance of Dominated Cluster Deletion, $C_{0}$ can be dominated by a set $D_{0}$ of size at most $d$. Then $D_{0}\cup B$ is dominating set of $G$ of size at most $q+d$.
@@ 4.1 Dominated Cluster Deletion
Note that $S$ may be the empty set in the third item above if $G-S^{\prime}$ has only one connected component. Figure˜3 shows examples of skeletons for dominated cluster deletion.
@@
Assume that Case 1 is not satisfied so that no vertices from the skeleton are in the dominating set $X$. Denote by $C^{\prime}_{0}$ the large connected component of $G-S^{\prime}$ for a set $S^{\prime}\supseteq S$ witnessing that $S$ is a skeleton, hence, $S$ contains exactly the vertices of $S^{\prime}$ that have a neighbor in $C^{\prime}_{0}$ and a second neighbor in $G-N[C^{\prime}_{0}]$. Let $C_{0}$ be the large connected component of GSG-S. Note that it contains $C^{\prime}_{0}$.
Let $A=S\cup C_{0}$ and $B=S\cup\bigcup_{1\leq j\leq m}C_{j}$, where $C_{1},\ldots,C_{m}$ are the other connected components of GSG-S. Then $(A,B)$ is a separation of order at most $|S|\leq k$, hence $|B|\leq q$, as $G$ is $(q,k)$-unbreakable. Pick an arbitrary element $v\in S$. By definition of $S$, $v$ has a neighbor in $C_{0}$ and a neighbor $w$ in a second connected component, say $C_{1}$ of GSG-S. Hence, $w$ must be a vertex of degree at most $q$ as its neighborhood lies in $G-C_{0}\subseteq B$. If $w$ is in $X$, then this vertex $v$ satisfies Case 2. Otherwise, $w$ must have a neighbor $x$ in $X$ (since $X$ is a dominating set). Since we are not in Case 1, this $x$ is not in $S$, and is therefore in $C_{1}$. Hence, similarly to $w$, we have that $x$ has degree at most $q$, and we are in Case 3.
Lemma˜4.6 gives rise to a bounded search tree branching algorithm that branches over the possible choices for a vertex from $S$ given by the conditions of the lemma.
@@
The skeleton-algorithm with parameter $q,k,d$ over a graph $G$ returns a family of sets of vertices, each set of size at most $k$. First, it computes a dominating set $X$ of size at most $q+d$ (returning the empty family if there is no such set). Second, it computes the set $X\cup Y\cup Z$, where $Y$ is the set of all vertices in the neighborhood $N(x)$ of a vertex $x\in X$ with $\deg(x)\leq q$, and the set $Z$ is the set of all vertices in the neighborhood $N(y)$ of a vertex $y\in Y$ where $\deg(y)\leq q$. Third, call the algorithm with parameters q,k1,dq,k-1,d on GvG-v for every $v$ in $X\cup Y\cup Z$ (and get a family $\mathcal{F}_{v}$). Finally, for every $v$ in $X\cup Y\cup Z$ add $v$ to every set in $\mathcal{F}_{v}$, and return the empty set and the union of all the families $\mathcal{F}_{v}$.
@@
Note that in the above definition, if a family $\mathcal{F}_{v}$ is empty, adding $v$ to every set still results in an empty family (e.g. in the case where GvG-v does not have a $q+d$ dominating set). Do not be confused with $\mathcal{F}_{v}$ having the empty set as one of the family members (e.g. for branches of the algorithm that stops before reaching $k=0$).
We now show that the skeleton algorithm 1) runs efficiently and outputs a small family of sets which is proved in Lemma˜4.9, and 2) if the given graph was a positive instance to Dominated Cluster Deletion, then the skeleton algorithm outputs all possible skeletons; see Lemma˜4.11.
@@
Let $t(\cdot)$ be a graph parameter satisfying ˜1, then the skeleton-algorithm runs in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$, and outputs a family of at most $g(q,k,d)$ sets where $f$ and $g$ are computable functions.
@@
Computing $X,Y$ and $Z$ takes time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$ thanks to the fact that $t(\cdot)$ satisfies ˜1. We then branch over $(q+d)q^{2}$ many choices for $v$, and the recursion has depth at most $k$.
@@
By induction on $k$. If $k=0$, we simply output the empty set. Otherwise, by Lemma˜4.3 there is such a set $X$, and by Lemma˜4.6 at least one element $v$ of $X$ is in $S$. For the correct choices of $v$, GvG-v is a positive instance for Dominated Cluster Deletion with parameter k1,dk-1,d, and by Lemma˜4.1, GvG-v is (q,k1)(q,k-1)-unbreakable, so the recursive call to Skeleton produces the expected sets by induction.
@@
Given a $(q,k)$-unbreakable graph $G$ and a set $S$ of at most $k$ vertices, we can test in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$, whether $S$ is a skeleton for Dominated Cluster Deletion with parameters $k$ and $d$ assuming $t(\cdot)$ satisfies ˜1.
@@
We need to check the existence of a set $S^{\prime}\supseteq S$ with $|S^{\prime}|\leq k$ such that:
@@
Note that by these conditions the large connected component $C_{0}$ of GSG-S is separated from the small connected components by $S$. Let $s=|S|$, what remains to be done is hence the following.
@@
    Find an optimal solution for $G-S-C_{0}$. This can be done in time $f(k,d)$, because this subgraph has size at most $q$ and yields a value $k^{\prime}$ of vertices that need to be deleted. If $k^{\prime}>k-s$, we conclude that this $S$ is not a skeleton because there is no possible $S^{\prime}$.
@@
    Test whether an additional set $W$ of size $k^{\prime\prime}=k-s-k^{\prime}$ can be deleted from $C_{0}$ such that $G[C_{0}-W]$ can be dominated by $d$ vertices. Note that by assumption on $S$ and $S^{\prime}$ the deletion of $W$ will not cause $C_{0}$ to break into multiple further connected components. Hence, we simply check whether $C_{0}$ is a positive instance of Annotated Partial Domination with parameter $k^{\prime\prime}$, where $R=B=C_{0}$, and $F=\emptyset$. If $S$ is a skeleton, such a set exists (as $S^{\prime}\cap C_{0}$ is a candidate) and hence, if $G$ is a positive instance of Dominated Cluster Deletion a positive solution will be found. Since $t(\cdot)$ is a graph parameter satisfying ˜1, we can test the existence of such $W$ in time $f(k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$. If no solution is found, then this $S$ is not a skeleton.
The combination of Lemmas˜4.6, 4.9, 4.11 and 4.13 yields the following theorem.
@@
Dominated Cluster Deletion on $(q,k)$-unbreakable graphs can be solved in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$, where $t(\cdot)$ is a parameter satisfying ˜1.
@@ 4.2 Elimination Distance to Dominated Clusters
First, let us properly define Elimination Distance to Dominated Clusters.
@@
  • $G$ is connected and there is a vertex $v$ such that GvG-v has elimination distance k1k-1 to $\mathscr{C}$, or
@@
In our case we measure, for every $d$, the elimination distance to the class $\mathscr{C}_{d}$ of $d$-dominated clusters, i.e. graphs where every connected component has a $d$-dominating set. We interchangeably write “elimination distance $k$ to $d$-dominated cluster”, and “a positive instance for $(k,d)$-Elimination Distance to Dominated Clusters”. For algorithmic purposes, the above definition is not the most useful. We will instead look at the set of vertices that needs to be deleted to get $d$-dominated clusters. While this set is of unbounded size, it has a nice tree structure that is much more useful for our algorithms. This characterization is pretty standard, see the definition of elimination order [10, Proposition 4.3]
@@
Let $G$ be a graph and $S\subseteq V(G)$. We say that $S$ is tree-structured if there is a rooted tree $T$ and a bijection $\lambda:S\rightarrow V(T)$ such that every path in $G$ between two vertices $u,v$ of $S$ must contain a common ancestor in $T$ of $\lambda(u)$ and $\lambda(v)$. The tree $T$ is called an elimination tree for $S$. The elimination depth of a tree-structured set with respect to the tree $T$ is the depth of $T$.The elimination depth of $S$ is the minimum elimination depth of $S$ over all elimination trees for $S$.
@@
A graph $G$ is a positive instance for $(k,d)$-Elimination Distance to Dominated Clusters if and only if there is a tree-structured set $S$ with elimination depth $k$ such that every connected component of GSG-S has a $d$-dominating set.
@@
Let $G$ be a $(q,k)$-unbreakable graph with at least $3q(k+q)$ vertices. Let $S$ be a tree-structured set with an elimination tree $T$ of depth at most $k$. Then
@@
    the largest connected component $C_{0}$ of GSG-S is uniquely determined, and
@@
    for every component $C$ of GSG-S there is a subset $S(C)\subseteq S$ of size at most $k$ that separates $C$ from the rest, that is, GS(C)G-S(C) has $C$ as one of its components.
@@
Let $C$ be a connected component of GSG-S. As $S$ is tree-structured by $T$, $C$ is not adjacent to two $\preccurlyeq_{T}$-incomparable elements, that is, the set $S(C)$ of neighbors of $C$ in $S$ consists of $\preccurlyeq_{T}$-comparable elements. $C$ is a connected component both of GSG-S and of GS(C)G-S(C), as all neighbors of $C$ in $S$ are collected in $S(C)$.
Assume now that there are two connected components $C_{0},C_{1}$ of GSG-S of size greater than $q$. However, $C_{0}$ and $C_{1}$ are separated by $S(C_{0})$ of size at most $k$, which contradicts the $(q,k)$-unbreakability of $G$. So there can only be one connected component of size greater than $q$.
We now show that there must be at least one such component. First note that there cannot be more than $3q$ connected component in GSG-S. Otherwise, there would be a balance separator of $T$ of size $1$, by deleting this vertex and its ancestors, we delete at most $k$ vertices and obtain a separation with both sides containing at least $q$ components of GSG-S, contradicting the unbreakability of $G$. Last, as there are only $3q$ components $C$ and $|S(C)|\leq k$ we have that $|S|\leq 3qk$ hence $|G-S|\geq 3q^{2}$ so on of the component must contain at least $q$ vertices.
Lemma 4.21 (Analog of Lemma˜4.3).
@@
The proof can be taken verbatim from Lemma˜4.3 with $S^{\prime}$ replaced by $S^{\prime}(C_{0})$.
@@
The following observation follows immediately from Lemma˜4.19. {observation} Every skeleton has size at most $k$.
Figure 4: Example of a skeleton for elimination distance to dominated clusters for parameter $k=3$ and $d=1$ on a (12,3)-unbreakable graph (left). The skeleton $S$ is in red, the witness $S^{\prime}$ also contains the crossed-out vertices. A tree structure for $S^{\prime}$ of elimination depth 33 is on the right.
Now the analog of Lemma˜4.6 also holds for skeletons for elimination distance to dominated clusters. The proof can be taken verbatim with $S^{\prime}$ replaced by $S^{\prime}(C_{0})$.
@@
Hence, the algorithm to guess a skeleton $S$ can be carried out in exactly the same way as for Dominated Cluster Deletion. We then guess a tree structure on $S$ (which takes time at most $\mathcal{O}(k!)$). Note that these vertices being adjacent to the large connected component, the tree structure is a linear order. In the final step, when we aim to extend a fixed skeleton $S$ to a tree-structured solution $S^{\prime}$, the only difference is that we have to solve the base problem Elimination Distance to Dominated Clusters for the small part $G-S-C^{\prime}_{0}$, which again can be simply solved by brute-force. Note that because $G-S-C^{\prime}_{0}$ is small, we can also find an elimination tree $T$ for $S^{\prime}$ and embed $S$ according to the originally guessed order. We conclude the main theorem of this section. The full algorithm for the annotated version is presented in the appendix.
@@
Elimination Distance to Dominated Clusters on $(q,k)$-unbreakable graphs can be solved in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$, where $t(\cdot)$ is a parameter satisfying ˜1.